uiz Instructions Question 5 The following boxplots show test scores for a first exam given in two different classes. Grades 100 90 80 70 60 class 1 class2 Variable Which class had the largest IQR? ® class 2 O class 1 Previous
Answer #1

Interquartile range (IQR) = 3rd quartile - 1st quartile

In approximate

IQR for class 1= 82-70 =12

IQR for calss 2= 80-62= 18

Therefore, the calss 2 has the largest IQR.